find three consecutive even integers if the third is 8 less than twice the second!


I need it today ! Please Help! CAn you also show me the work please ? Thank you for who ever answers it

Answer by scott8148(6628)

n , n+2 , n+4

(n + 4) + 8 = 2(n + 2)

n + 12 = 2n + 4